OpenStudy (anonymous):

simplify ((c/d)-(d/c))/((1/d)-(1/c))

OpenStudy (anonymous):

\[\frac{\frac{c^2-d^2}{cd}}{\frac{c-d}{dc}}\]

OpenStudy (anonymous):

\[\frac{c^2-d^2}{cd}*\frac{cd}{c-d}=?\] <__ Can you finish it now?